OpenStudy (anonymous):

the perimeter of a rectangle is 532 yards. the length is 9 yards less than the quadruple the width. What is the length and width

OpenStudy (hockeychick23):

Length=4w-9 Width= w There's 4 sides to a rectangle so: 4w-9+4w-9+w+w= 532

OpenStudy (anonymous):

@hockeychick23 so what are the individual dimensions?

OpenStudy (hockeychick23):

10w-18=532 10w-18+18=532+18 > 10w= 550 > w= 55, L= 211